# MLAAD: The Multi-Language Audio Anti-Spoofing Dataset

### v7

@@ Abstract
Text-to-Speech (TTS) technology offers notable benefits, such as providing a voice for individuals with speech impairments, but it also facilitates the creation of audio deepfakes and spoofing attacks. AI-based detection methods can help mitigate these risks; however, the performance of such models is inherently dependent on the quality and diversity of their training data. Presently, the available datasets are heavily skewed towards English and Chinese audio, which limits the global applicability of these anti-spoofing systems. To address this limitation, this paper presents the Multi-Language Audio Anti-Spoofing Dataset (MLAAD), version 7, created using 101 TTS models, comprising 52 different architectures, to generate 485.3 hours of synthetic voice in 40 different languages.
@@ I Introduction
Release Date Hours Models Languages Distributed via
v1 17.01.2024
& 21 22 zip v2 28.02.2024 63.0 21 22 zip v3 16.04.2024 163.9 54 23 zip v4 24.09.2024 175.0 59 23 zip v5 02.11.2024 398.3 87 38 zip + Huggingface v6 26.04.2025 420.7 91 38 Huggingface v7 11.07.2025 485.3 101 40 Huggingface
TABLE I: Overview of MLAAD dataset versions, including release dates, total duration of speech data, number of TTS models, supported languages, and distribution method.
@@
To address these issues, we introduce and continuosly expand our MLAAD dataset (c.f. Section I), a large-scale corpus of state-of-the-art audio fakes encompassing 40 languages. This dataset, based on the M-AILABS Speech Dataset [10], includes 485.3 hours of synthesized speech, created by 101 state-of-the-art TTS models. By training state-of-the-art deepfake detection models using MLAAD, we show that the resulting models achieve high performance scores in cross-dataset evaluations, indicating practicality in real-world conditions.
@@ III-B Formatting
The outputs are organized into directories describing language and model type. Each directory contains the synthesized audio files and an accompanying “meta.csv” file, which encapsulates various data fields as indicated in  Table III. The “path” field specifies the location of the synthesized audio file, relative to the base path of MLAAD. The “original_file” field denotes the filepath of the baseline audio, relative to the base path of the M-AILABS dataset. The “language” field identifies the language of the audio file, while “is_original_language” indicates if the language was originally included in M-AILABS, or if the transcript has been translated. The “duration” field shows the length of the audio file in seconds, and “training_data” reveals the dataset used to train the TTS model. The “model_name” field provides a unique identifier for the model, such as “facebook/mms-tts-eng” for the Facebook Massively Multilingual Speech (MMS) TTS model [61], while “architecture” describes the TTS model architecture (such as VITS). Lastly, “transcript” represents the textual transcript of the audio file.
@@ III-C Statistics
The dataset encompasses 485.3 hours of synthesized speech across 40 languages, generated using 101 state-of-the-art TTS models, comprising 52 different architectures. We distribute only these synthesized samples, for the originals please refer to the M-AILABS dataset.
@@ V Conclusion
In this paper, we present a new audio dataset, which comprises voice spoofs in 40 languages by 101 TTS models, composed of 52 architectures. When evaluating MLAAD’s usefulness as a training resource, we observe that MLAAD complements existing research. It consistently outperforms related work such as InTheWild or FakeOrReal with respect to the trained models’ cross-dataset generalization capability. Furthermore, in comparison with the renowned ASVspoof 2019 dataset, MLAAD proves to be a complementary resource. Across eight datasets, MLAAD and ASVspoof 2019 alternately outperformed each other, both excelling on four datasets: Training the SSL W2V2 model on ASVspoof19 results in the highest performance on ASVspoof21-DF, FakeOrReal, InTheWild, and MLAAD. Similarly, training SSL W2V2 or WhisperDF on MLAAD leads to optimal performance on ASVspoof19, ASVspoof21-LA, Voc.v, and WaveFake. We note that an evaluation of the multi-lingual anti-spoofing capabilities has to remain a topic for future work since there are no other comparable multi-lingual datasets against which to evaluate MLAAD. However, the strong results on WaveFake, which comprises both English and Japanese, suggest that the learning models profit from the diversity of languages in MLAAD.

### v8

@@ Abstract
Text-to-Speech (TTS) technology offers notable benefits, such as providing a voice for individuals with speech impairments, but it also facilitates the creation of audio deepfakes and spoofing attacks. AI-based detection methods can help mitigate these risks; however, the performance of such models is inherently dependent on the quality and diversity of their training data. Presently, the available datasets are heavily skewed towards English and Chinese audio, which limits the global applicability of these anti-spoofing systems. To address this limitation, this paper presents the Multi-Language Audio Anti-Spoofing Dataset (MLAAD), version 8, created using 119 TTS models, comprising 58 different architectures, to generate 570.3 hours of synthetic voice in 40 different languages.
@@ I Introduction
Release Date Hours Models Languages Distributed via
v1 17.01.2024
21 22 zip v2 28.02.2024 63.0 21 22 zip v3 16.04.2024 163.9 54 23 zip v4 24.09.2024 175.0 59 23 zip v5 02.11.2024 398.3 87 38 zip + Huggingface v6 26.04.2025 420.7 91 38 Huggingface v7 11.07.2025 485.3 101 40 Huggingface v8 01.10.2025 570.3 119 40 Huggingface
TABLE I: Overview of MLAAD dataset versions, including release dates, total duration of synthesized speech, number of TTS models, supported languages, and distribution method.
To address these issues, we introduce and continuosly expand our MLAAD dataset (c.f. Section I), a large-scale corpus of state-of-the-art audio fakes encompassing 40 languages. This dataset, based on the M-AILABS Speech Dataset [10], includes 570.3 hours of synthesized speech, created by 119 state-of-the-art TTS models. By training state-of-the-art deepfake detection models using MLAAD, we show that the resulting models achieve high performance scores in cross-dataset evaluations, indicating practicality in real-world conditions.
@@ III-B Formatting
The outputs are organized into directories describing language and model type. Each directory contains the synthesized audio files and an accompanying “meta.csv” file, which encapsulates various data fields as indicated in  Table III. The “path” field specifies the location of the synthesized audio file, relative to the base path of MLAAD. The “original_file” field denotes the filepath of the baseline audio, relative to the base path of the M-AILABS dataset. The “language” field identifies the language of the audio file, while “is_original_language” indicates if the language was originally included in M-AILABS, or if the transcript has been translated. The “duration” field shows the length of the audio file in seconds, and “training_data” reveals the dataset used to train the TTS model. The “model_name” field provides a unique identifier for the model, such as “facebook/mms-tts-eng” for the Facebook Massively Multilingual Speech (MMS) TTS model [61], while “architecture” describes the TTS model architecture (such as VITS), and “transcript” represents the textual transcript of the audio file. Lastly, “reference_speaker” specifies the reference speaker used to generate the audio file. Depending on the TTS model, this value is either a file path pointing to an M-AILABS audio file or a speaker name (for models that provide a fixed set of speakers). Note that this key was introduced only in Version “v8”, and may therefore be missing from previous “meta.csv” files.
@@ III-C Statistics
The dataset encompasses 570.3 hours of synthesized speech across 40 languages, generated using 119 state-of-the-art TTS models, comprising 58 different architectures. We distribute only these synthesized samples, for the originals please refer to the M-AILABS dataset.
@@ V Conclusion
In this paper, we present a new audio dataset, which comprises voice spoofs in 40 languages by 119 TTS models, composed of 58 architectures. When evaluating MLAAD’s usefulness as a training resource, we observe that MLAAD complements existing research. It consistently outperforms related work such as InTheWild or FakeOrReal with respect to the trained models’ cross-dataset generalization capability. Furthermore, in comparison with the renowned ASVspoof 2019 dataset, MLAAD proves to be a complementary resource. Across eight datasets, MLAAD and ASVspoof 2019 alternately outperformed each other, both excelling on four datasets: Training the SSL W2V2 model on ASVspoof19 results in the highest performance on ASVspoof21-DF, FakeOrReal, InTheWild, and MLAAD. Similarly, training SSL W2V2 or WhisperDF on MLAAD leads to optimal performance on ASVspoof19, ASVspoof21-LA, Voc.v, and WaveFake. We note that an evaluation of the multi-lingual anti-spoofing capabilities has to remain a topic for future work since there are no other comparable multi-lingual datasets against which to evaluate MLAAD. However, the strong results on WaveFake, which comprises both English and Japanese, suggest that the learning models profit from the diversity of languages in MLAAD.